Actions
Ops/Comm Scheduling meeting : 30/10/2017¶
Location : CR21 at 14:00 pm
Present: CM,KL,PK,MU,CW,HN,PF,DR,CR,SB,MT
Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517
- DC Update
- Shutdown activities - the list so far
Name | Task | Time required | Done? |
Henry | Restart conventional magnet controls crate | 1-2 hours (?) | Wednesday |
Josef | Replace filters in 8 compressors on mezzanine | 1-2 days | In progess |
Ajit | Test AMI controllers | 0.5 - 1 day | Arranging for Friday |
Mark | Partial warming of DS to clean vacuum space and recool | 10 days | Cooling - 4 days to go |
Ajit/MOM | Review current alarms | 1 day | Went through alarms with CM, all understood, some state changes req'd |
Joe/Henry | Change target patch fibres | 0.5 days | In progress, Joe down Tues/Weds. Inspect in synch |
John/Colin | Hall AC work? | ? | Maintenance this week |
- Others?
Tracker - Check David has finished his work. Need calibration run next week. Melissa & Durga to help wrap this up.
- Next user cycle : 2017/03 : Nov 14 - Dec 20
- Shifts are now open. Please sign up. 32/48 shifts have been assigned.
- MOMs will be : Durga, Victoria+Melissa, Paolo
- Run up will probably be on Nov 14th with data-taking starting on Nov 15th.
- AOB
-HN Reboot data mover
-Paolo add more disk space on MICEraid5
-Need to give Tony Millington heads up on survey requirements in Nov/Dec once the run plan has settled (3 Absorber changes)
-Ajit looking at producing automated cooling channel magnet ramp down procedure for use by MOM's
Updated by MacWaters, Craig almost 6 years ago · 5 revisions